OpenStudy (anonymous):

Describe the vertical asymptotes and holes for the graph of y=x-4/x^2+3x+2

OpenStudy (aum):

Step 1: Factor the denominator

OpenStudy (anonymous):

(x+2)(x+1)

OpenStudy (jdoe0001):

vertical asymptotes for a rational expression occur when the fraction becomes undefined, and that only happens when the denominator turns to 0 thus \(\bf y=\cfrac{x-4}{{\color{brown}{ x^2+3x+2}}}\qquad x^2+3x+2=0\implies y=\cfrac{x-4}{0} \\ \quad \\ \textit{thus the vertical asymptotes occur at }{\color{brown}{ x^2+3x+2}}=0\) and you're correct, are (x+2)(x+1)

OpenStudy (jdoe0001):

which means \(\large { x^2+3x+2=0\\ \quad \\\implies (x+2)(x+1)=0\to \begin{cases} x+2=0\to x=-2\\ x+1=0\to x=-1 \end{cases} \\ \quad \\ asymptotes\to \begin{cases} y=-2\\ y=-1 \end{cases} }\)

OpenStudy (jdoe0001):

and the "holes" occur if x = -2 or x = -1

OpenStudy (jdoe0001):

hmmmm wait a second... is a vertical asymptote... so the equation should be x-based... so \(\large {x^2+3x+2=0\\ \quad \\ (x+2)(x+1)=0\to \begin{cases} x+2=0\to x=-2\\ x+1=0\to x=-1 \end{cases} \\ \quad \\ asymptotes\ and \ holes\to \begin{cases} x=-2\\ x=-1 \end{cases}}\)

OpenStudy (anonymous):

So there are holes because x=-2 and x=-1?

OpenStudy (jdoe0001):

yes, because if we set x = -2 or -1 our denominator will turn to 0 and thus our equation will be undefined \(y=\cfrac{x-4}{{\color{brown}{ x^2+3x+2}}}\qquad x^2+3x+2=0\implies y=\cfrac{x-4}{0} \\ \quad \\ y=\cfrac{x-4}{{\color{brown}{ (x+2)(x+1)}}} \begin{cases} y=\cfrac{x-4}{{\color{brown}{ ({\color{blue}{ -2}}+2)(x+1)}}}\to y=\cfrac{x-4}{0(x+1)}\to y=\cfrac{x-4}{0} \\ \quad \\ y=\cfrac{x-4}{{\color{brown}{ (x+2)({\color{blue}{ -1}}+1)}}}\to y=\cfrac{x-4}{(x+2)0}\to y=\cfrac{x-4}{0} \end{cases} \)